OpenStudy (anonymous):

What is the value of a in the equation 5a – 10b = 45, when b = 3? 3 15 21 39

OpenStudy (anonymous):

5a - 10b = 45 Plug in 3 for 'b': 5a - 10(3) = 45 Can you multiply -10 * 3?

OpenStudy (anonymous):

i got -30

OpenStudy (anonymous):

Yes, so we have: 5a - 30 = 45 Add 30 to both sides, what's 45 + 30?

OpenStudy (anonymous):

75

OpenStudy (anonymous):

Yes, so we have: 5a = 75 Now divide 5 to both sides, what's 75 / 5?

OpenStudy (anonymous):

15! =D

OpenStudy (anonymous):

Yes, so a = 15!

OpenStudy (anonymous):

Thanks mann ' you very good at math man! =D

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

I need help doing a quiz i just got bro ' can I you help me through it?